OpenStudy (anonymous):

help! will medal! solve: x/5>-3 - a. x>15 - b. x>-15 - c. x<15 - d. x<-15 - Please dont give me the answer, walk me through the problem! < OR > means less than or equal to OR greater - - than or equal to

OpenStudy (igreen):

\(\dfrac{x}{5} \ge -3\) is this it? @pigglet

OpenStudy (gahm8684):

Let's solve your inequality step-by-step. x 5 >−3 Step 1: Simplify both sides of the inequality. 1 5 x>−3 Step 2: Divide both sides by 1/5. 1 5 x 1 5 > −3 1 5 x>−15 Answer: x>−15

OpenStudy (anonymous):

a is the answer

OpenStudy (anonymous):

yes, @iGreen

OpenStudy (igreen):

@gahm8684 @aero87 Don't give out answers.

OpenStudy (igreen):

\(\dfrac{x}{5} \ge -3\) Okay, to get x by itself we have to multiply 5 to both sides. So what's 5 * -3? @pigglet

OpenStudy (anonymous):

just multipy both sides by 5. that's all.

OpenStudy (igreen):

Your answer is incorrect. @aero87

OpenStudy (anonymous):

I clearly said, dont give me the answer, walk me through it @gahm8684 @aero87 @iGreen -15?

OpenStudy (igreen):

Yes, therefore that gives us: \(x \ge -15\)

OpenStudy (igreen):

5 was being divided by 'x'..so we have to do the opposite and multiply it.

OpenStudy (anonymous):

so, its b?

OpenStudy (igreen):

Yep.

OpenStudy (anonymous):

thanks!

OpenStudy (igreen):

No problem.